### v1

# Relative uniform convergence and Archimedean property in pre-ordered vector spaces

Abstract
It is proved that, for a pre-ordered vector space $X$, the quotient space $(X/A,[W])$ is the Archimedeanization of $X$, where $W$ is the closure of the positive wedge $X_{+}$ in the ru-topology, $A=W\cap(-W)$, and $[W]$ is the quotient set of $W$ in $X/A$.
Eduard Emelyanov1
11 Sobolev Institute of Mathematics, Novosibirsk, Russia
Keywords: pre-ordered vector space, relative uniform convergence, Archimedean wedge, Archimedeanization
MSC2020: 46A40, 47B60, 47B65
1 Introduction
Archimedean property is important in the functional representation of various analytic and algebraic structures. For example, Kadison’s representation theorem [7] tells that every ordered real vector space with an Archimedean order unit is order isomorphic to a vector subspace of the space of continuous real valued functions on a compact Hausdorff space. Kadison’s theorem inspired M.D. Choi and E.G. Effros to obtain an analogous representation theorem for self-adjoint subspaces of unital $C^{\ast}$-algebras that contain the unit [2].
There is a plenty of recent works by many mathematicians who studied the Archimedean property in connection to various applications in algebra, analysis and quantum physics. In the breakthrough paper [12], V.I. Paulsen and M. Tomforde have developed a theory of ordered *-vector spaces. In their theory, Archimedeanization of ordered vector spaces with order units plays a crucial role. The construction of Archimedeanization of ordered vector spaces with an order unit [12] was extended to arbitrary ordered vector spaces in [3]. We also mention one more important direction in analysis (based on representation theorems for Archimedian structures), the construction of free analytic spaces that was initiated by B. de Pagter and A.W. Wickstead in their pioneered work [11].
The present paper is devoted to an interplay between the Archimedean property and relatively uniform convergence in (pre-)ordered vector spaces. In Theorem 3.3, we prove that the quotient set $[W]$ of the $\tau_{ru}$-closure of the positive wedge $X_{+}$ of a pre-ordered vector space $X$ is an Archimedean cone in quotient space $X/(W\cap(-W))$. The main result of our paper is Theorem 3.5 which establishes that $(X/(W\cap(-W)),[W])$ is the Archimedeanization of $X$.
2 Preliminaries
In what follows, vector spaces are real and operators are linear. We recall few well known definition. For further unexplained terminology and notations we refer to [1, 5, 8].
A subset $W$ of a vector space $X$ is a wedge if $W+W\subseteq W\ \ \text{\rm and}\ \ tW\subseteq W\ \ \text{\rm for}\ \ t\geqslant 0$. A wedge $C$ is a cone whenever $C\cap(-C)=\{0\}$. Under an ordered vector space (briefly, OVS) we understand a vector space $X$ together with a cone (denoted by $X_{+}$ and called a positive cone of $X$). Replacing a cone by a wedge, we also consider a pre-ordered vector space (briefly, POVS). A wedge $W$ in a vector space $X$ is majorizing (or generating) if
[…]
for every $x\in X$.
Every OVS (POVS) $X$ is equipped with a partial (pre-)order
$x\leqslant y\Longleftrightarrow y-x\in X_{+}.$
Every pair of vectors $\{a,b\}$ in an OVS (POVS) $X$ produces an order interval (possibly empty, if $a\not\leqslant b$):
$a,b]:=\{x\in X:a\leqslant x\leqslant b\}=(a+X_{+})\cap(b-X_{+}).$
A subset $A$ of a POVS is called an order ideal in $X$, whenever $A$ is a vector subspace of $X$ such that $a,b\in A\Longrightarrow[a,b]\subseteq A$.
An operator $T:X\to Y$ between POVSs is positive (order bounded) if $T$ takes $X_{+}$ into $Y_{+}$ (order intervals $X$ into order intervals).
A wedge $X_{+}$ in a POVS $X$ is (or, $X$ is):
  1. -
    Archimedean if $y\leqslant 0$ whenever $ny\leqslant u\in X_{+}$ for all $n\in\mathbb{N}$.
  2. -
    almost Archimedean if $\bigcap\limits_{n\in\mathbb{N}}\bigl[-\frac{1}{n}u,\frac{1}{n}u\bigl]=\{0\}$ for every $u\in X_{+}$.
Clearly, each almost Archimedean POVS must be an OVS. Moreover, every Archimedean OVS is almost Archimedean. The converse is not true even for two-dimensional OVSs (see Example 2.5 in [3]). It is also straightforward to see: a POVS $X$ is almost Archimedean iff $X_{+}$ does not contain a straight line.
A net $(x_{\alpha})$ in a POVS $X$ ru-converges to $x\in X$ (shortly, $x_{\alpha}\operatorname{\xrightarrow[]{{\it ru}}}x$) if, for some (a regulator of the convergence) $w\in X_{+}$, there exists a sequence $(\alpha_{n})$ of indices with $x_{\alpha}-x\leqslant\pm\frac{1}{n}w$ for $\alpha\geqslant\alpha_{n}$. Whenever we need to specify a regulator $w$, we write $x_{\alpha}\operatorname{\xrightarrow[]{{\it ru}}}x(w)$.
Clearly, every ru-convergent net in an OVS $X$ has a unique ru-limit iff $X$ is almost Archimedean. As the ru-convergence is sequential, we can always restrict ourselves to ru-convergent sequences. The ru-convergence is an abstraction of the classical uniform convergence of functions in $C[0,1]$. The notion of ru-convergence was introduced by L.V. Kantorovich [9].
3 Closedness of the positive wedge in ru-topology vs Archimedean property in pre-ordered vector spaces
Let $X/A$ be a quotient space. A quotient set of $B\subseteq X$ is the set $B]=\{[x]\in X/A:x\in B\}$ of all equivalence classes formed from a set $B$ in the quotient space $X/A$.
The following lemma is a standard fact.
Lemma 3.1.
Let $W$ be a wedge in a vector space $X$. Then $[W]$ is a cone in $X/A$, where $A=W\cap(-W)$. Moreover, $[W]$ is majorizing in $X/A$ whenever $W$ is majorizing in $X$.
Proof.
Clearly, $A$ is a subspace of $X$ and $[W]$ is a wedge in $X/A$. To see that $[W]$ is a cone it suffices to show $W]\cap-[W]\subseteq A$. Let $x]\in[W]\cap-[W]$. Then $x=x_{1}+A=-x_{2}+A$ for some $x_{1},x_{2}\in W$. It follows $x\in(W+W)\cap(-W-W)\subseteq W\cap(-W)=A$, and hence $[x]=0$. The rest is trivial. ∎
An OVS $X$ is a vector lattice if $x^{+}:=\inf\bigl\{X_{+}\cap(x+X_{+})\bigl\}$ exists for every $x\in X$. It is easy to see that every almost Archimedean vector lattice is Archimedean. An order ideal $A$ in a vector lattice $X$ will be referred to as a lattice ideal whenever it satisfies the property: $a\in A\Longrightarrow|a|\in A$.
W.A.J. Luxemburg and L.C-Jr. Moore proved in [10] that the ru-closed subsets of a vector lattice $X$ are exactly the closed sets of the so-called ru-topology on $X$. This topology can be also introduced on an arbitrary POVS (cf., [8]).
Definition 3.1.
The ru-topology, denoted by $\tau_{ru}$, on a POVS $X$ is determined as follows: a subset $S$ of $X$ is $\tau_{ru}$-closed whenever $S\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ implies $x\in S$. The $\tau_{ru}$-closure of a subset $S$ of $X$ is denoted by $\overline{S}_{ru}$.
It follows immediately from Definition 3.1 that the ru-topology on a POVS $X$ is the strongest topology $\tau$ on $X$ with the property $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$.
The ru-topology is not linear in general. It is proved in [10, Theorem 4.2] that every lattice homomorphism between vector lattices is continuous in the ru-topology. Indeed, it is true for arbitrary positive operators between POVSs, and if additionally the positive wedge in the co-domain is majorizing, it is also true for every order bounded operator.
Lemma 3.2.
Let $T:X\to Y$ an operator between POVSs $X$ and $Y$.
  1. $i)$
    If $T$ is positive then $T$ is $\tau_{ru}$-continuous.
  2. ii)ii)
    If $Y_{+}$ is majorizing and $T$ is order bounded then $T$ is $\tau_{ru}$-continuous.
Proof.
Let $S\subseteq Y$ $\tau_{ru}$-closed. It suffices to show that $T^{-1}(S)$ is $\tau_{ru}$-closed. To this end, let $T^{-1}(S)\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$. By passing to a subsequence, we may suppose $x_{n}-x\leqslant\pm\frac{1}{n}u$ for some $u\in X_{+}$ and all $n\in\mathbb{N}$. By the definition of $\tau_{ru}$-topology, it remains to show $x\in T^{-1}(S)$.
If $T\geqslant 0$ then $Tx_{n}-Tx\leqslant\pm\frac{1}{n}Tu$, or $Tx_{n}\operatorname{\xrightarrow[]{{\it ru}}}Tx$. Therefore, $S\ni Tx_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}Tx$ and, since $S$ is $\tau_{ru}$-closed, we infer $Tx\in S$, and hence $x\in T^{-1}(S)$.
If $Y_{+}$ is majorizing and $T$ is order bounded, then $T$ is ru-continuous by [5, Theorem 2.4 ii)] As $T^{-1}(S)\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ then $S\ni Tx_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}Tx$. Thus, $Tx\in S$, and again $x\in T^{-1}(S)$. ∎
It was observed by T. Ito [6] that a vector lattice $X$ is Archimedean iff $X_{+}$ is ru-closed. We show that it is happened also with every POVS.
Lemma 3.3.
A POVS $X$ with a positive wedge $X_{+}$ is Archimedean iff $X_{+}$ is ru-closed.
Proof.
$\Longrightarrow$: Let $X$ be Archimedean and $X_{+}\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x_{0}(w)$. By passing to a subsequence, we may suppose $\frac{1}{n}w\leqslant x_{0}-x_{n}\leqslant\frac{1}{n}w$ for all $n\in\mathbb{N}$. Then, $x_{0}=(x_{0}-x_{n})+x_{n}\geqslant x_{0}-x_{n}\geqslant-\frac{1}{n}w$, and hence $w\geqslant n(-x_{0})$ for all $n\in\mathbb{N}$. Since $X$ is Archimedean, $(-x_{0})\leqslant 0$, or $x_{0}\in X_{+}$.
[…]
Lemma 3.3 should be compared with the following classical fact (cf., [1, Lemma 2.3]).
Proposition 3.1.
If a POVS $X$ admits a linear topology $\tau$ for which its wedge $X_{+}$ is $\tau$-closed, then $X$ is Archimedean.
The $\tau_{ru}$-topology is not necessarily linear. However, as it was mentioned above, the $\tau_{ru}$-topology on a POVS $X$ is the strongest topology $\tau$ on $X$ such that $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$. Therefore, we obtain the following corollary of Lemma 3.3.
Corollary 3.1.
A POVS $X$ is Archimedean iff $X_{+}$ is $\tau$-closed in some topology $\tau$ on $X$ possessing the property: $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$.
One of the most exciting results concerning interrelations between Archimedean cones and ru-convergence is the Luxemburg-Moore-Veksler theorem:
Theorem 3.2.
([10, 13]) Let $A$ be a lattice ideal in a vector lattice $X$. Then
[…]
The forward implication of Theorem 3.2 is almost trivial in the OVS case [3, Proposition 3.2(a)]. We do not have yet an extension of Theorem 3.2 to any reasonable class of POVS beside vector lattices. However, for a very natural ru-closed order ideal $A$ in an arbitrary POVS $X$ the quotient $X/A$ turns to be Archimedean.
Theorem 3.3.
Let $X$ be a POVS with a positive wedge $X_{+}$. Then $[W]$ is an Archimedean cone in $X/A$, where $W$ is the $\tau_{ru}$-closure of $X_{+}$ and $A=W\cap(-W)$. Moreover, the cone $[W]$ is majorizing whenever $X_{+}$ is majorizing.
Proof.
Clearly, $A$ is an order ideal in $X$. By Lemma 3.1, $[W]$ is a cone in $X/A$. The set $W$ (and, also W-W) is $\tau_{ru}$-closed. Let $n[x]\leqslant[y]$ for all $n\in\mathbb{N}$ and some $x\in X,y\in W$. Then $x]\leqslant\frac{1}{n}[y]$, and hence $x-\frac{1}{n}y]\in-[W]$ for all $n\in\mathbb{N}$. So, there exists $(a_{n})_{n}$ in $A$ such that $x-\frac{1}{n}y+a_{n}\in-W$ for all $n\in\mathbb{N}$. Note that $a_{n}\operatorname{\xrightarrow[]{{\it ru}}}0$ (with a regulator 0). As W-W is $\tau_{ru}$-closed, we infer from $\frac{1}{n}y\operatorname{\xrightarrow[]{{\it ru}}}0$ and $a_{n}\operatorname{\xrightarrow[]{{\it ru}}}0$ that $x-\frac{1}{n}y+a_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}x$, and hence $x\in-W$. It follows $x]\in-[W]$, or $x]\leqslant 0$. So, $(X/A,[W])$ is Archimedean. Due to Lemma 3.1, $[W]$ is majorizing whenever $X_{+}\subseteq W$ is majorizing. ∎
It was proved in [12] that every OVS with an order unit has a unique (up to an order isomorphism) Archimedization. This result was extend further to arbitrary OVS in [3]. The following definition was introduced in [3, Definition 5].
Definition 3.2.
Let $X$ be a POVS. Consider the category $\mathcal{C}_{Arch}(X)$ whose objects are pairs $\left\langle Y,\phi\right\rangle$, where $Y$ is an Archimedean OVS and $\phi:X\to Y$ is a positive operator, and morphisms $\left\langle Y_{1},\phi_{1}\right\rangle\to\left\langle Y_{2},\phi_{2}\right\rangle$ are positive operators $T_{12}:Y_{1}\to Y_{2}$ with $T_{12}\circ\phi_{1}=\phi_{2}$. If $\left\langle Y_{0},\phi_{0}\right\rangle$ is an initial object of $\mathcal{C}_{Arch}(X)$, then the OVS $Y_{0}$ is called the Archimedeanization of $X$.
By [3, Corollary 3.5], any almost Archimedean cone can be embedded into an Archimedean cone. In particular, the Archimedeanization of any almost Archimedean OVS space $X$ can be obtained via replacing $X_{+}$ by the intersection of all Archimedean cones containing $X_{+}$. The following result is proved in [3, Theorem 1].
Proposition 3.4.
Every OVS has the Archimedeanization.
We present a construction of the Archimedeanization (based on Theorem 3.3), which seems to be more elementary than one given in [3].
Theorem 3.5.
Let $X$ be a POVS with a positive wedge $X_{+}$. Let $W$ is the $\tau_{ru}$-closure of $X_{+}$ and $A=W\cap(-W)$. Then $(X/A,[W])$ is the Archimedeanization of $X$.
Proof.
A quotient map $\phi_{0}:X\to(X/A,[W])$ is positive. By Theorem 3.3, $\left\langle(X/A,[W]),\phi_{0}\right\rangle$ is an object of $\mathcal{C}_{Arch}(X)$. We need to show that $\left\langle(X/A,[W]),\phi_{0}\right\rangle$ is initial. To this end, take an arbitrary object $\left\langle Y,\phi\right\rangle$ of $\mathcal{C}_{Arch}(X)$. It remains to find a positive operator $T:(X/A,[W])\to Y$ such that $T\circ\phi_{0}=\phi$. Since $Y$ is Archimedean, it follows from Lemma 3.3 that $Y_{+}$ is ru-closed, and hence is $\tau_{ru}$-closed. First, we show $\phi(W)\subseteq Y_{+}$. Let $x\in W$. There exists a net $(x_{\alpha})$ such that $X_{+}\ni x_{\alpha}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}x$. By Lemma 3.2, the operator $\phi:X\to Y$ is $\tau_{ru}$-continuous. Therefore, we obtain $Y_{+}\ni\phi(x_{\alpha})\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}\phi(x)$. As $Y_{+}$ is $\tau_{ru}$-closed, $\phi(x)\in Y_{+}$. Thus, $\phi(W)\subseteq Y_{+}$ and $\phi(-W)\subseteq-Y_{+}$, and hence $\phi(A)\subseteq Y_{+}\bigcap(-Y_{+})=\{0\}$. So, $T:(X/A,[W])\to Y$ such that $T[x]:=\phi(x)$ is well defined and positive. Clearly, $T\circ\phi_{0}=\phi$. ∎
For applications of Theorem 3.5, it is useful to describe constructively the $\tau_{ru}$-closure of the positive wedge $X_{+}$ of a POVS $X$. The idea comes from [10], where the pseudo uniform closure of a subset of a vector lattice was studied.
Adopting the idea to the POVS setting, for a subset $S$ of a POVS $X$ we denote by $S^{\prime}_{ru}$ the set of all $x\in X$ such that $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ for some sequence $(x_{n})_{n}$ in $S$. Denote by $S^{(\alpha)}_{ru}$ the set $S^{(\alpha-1)}_{ru}$ when the ordinal $\alpha$ is not limit, otherwise $S^{(\alpha)}_{ru}:=\bigcup\limits_{\beta\in\alpha}S^{(\beta)}_{ru}$. It is straightforward to see that $S^{(\alpha_{1})}_{ru}\subseteq S^{(\alpha_{2})}_{ru}$ whenever $\alpha_{1}\leqslant\alpha_{2}$, and $\overline{S}_{ru}=\bigcup\limits_{\alpha\in\omega_{1}}S^{(\alpha)}_{ru}$, where $\omega_{1}$ is the first uncountable ordinal (cf., [10, Theorem 3.3] in the vector lattice setting). In particular, we have $\overline{S}_{ru}=S^{\prime}_{ru}\ \Longleftrightarrow\ S^{\prime}_{ru}=(S^{\prime}_{ru})^{\prime}_{ru}$. So, in the notations of Theorem 3.5, $W=\overline{(X_{+})}_{ru}=\bigcup\limits_{\alpha\in\omega_{1}}(X_{+})^{(\alpha)}_{ru}$. Note that each $(X_{+})^{(\alpha)}_{ru}$ is a wedge in $X$.
The statement of Theorem 3.5 motivates the following notion.
Definition 3.3.
A POVS $X$ has the Archimedeanization type $\alpha(X)$, whenever $\alpha(X)$ is the first ordinal $\alpha$ such that $\overline{(X_{+})}_{ru}=(X_{+})^{(\alpha)}_{ru}$.
It can be seen easily:
  1. -
    $\alpha(X)\leqslant 1$ for every almost Archimedean POVS $X$, and $\alpha(X)=0$ iff $X$ is Archimedean.
  2. -
    $0\leqslant\alpha(X)<\omega_{1}$ for every POVS $X$, and $\alpha(X)=2$ for the vector lattice $X$ in Nakayama’s example (cf., Example 6.4 in [10]).
Let us consider the following order ideals in a POVS $X$: the set $I_{0}(X):=\{0\}$ of all infinitesimals of $X$ of order 0; the set:
$I_{1}(X)=I(X):=\bigl\{x\in X:(\exists y\in X)(\forall n\in\mathbb{N})\pm x\leqslant\frac{1}{n}y\bigl\}$
of all infinitesimals of order 1 (or, just infinitesimals of $X$); and the set
$I_{\alpha}(X):=\bigl\{x\in X:[x]\in I\bigl(X/_{\bigcup_{\gamma\in\alpha}I_{\gamma}(X)}\bigl)\bigl\}.$
of all infinitesimals of order $\alpha$, where $\alpha$ is an arbitrary ordinal >1>1.
It can be seen that $I_{\alpha_{1}}(X)\subseteq I_{\alpha_{2}}(X)\subseteq(X_{+})^{(\alpha_{2})}_{ru}$ for $\alpha_{1}\leqslant\alpha_{2}$. Clearly, a POVS $X$ is almost Archimedean iff $I(X)=\{0\}$. If a POVS $X$ has a nonzero order unit $u\in X$ (i.e., $u$ satisfies $X=\bigcup\limits_{n\in\mathbb{N}}[-nu,nu]$) then $u\notin I(X)$.
Definition 3.4.
By the almost Archimedeanization type $\lambda(X)$ (cf., [3, 4]) of a POVS $X$ we understand the first ordinal $\lambda$ such that $I_{\lambda+1}(X)=I_{\lambda}(X)$.
It can be shown easily:
  1. -
    $\lambda(X)=0$ iff $X$ is almost Archimedean, in particular $\lambda(X)=0$ for every normed lattice.
  2. -
    $I\bigl(\ell^{\infty}/c_{00}\bigl)=I\bigl(c_{0}/c_{00}\bigl)=c_{0}/c_{00}$, and hence $I_{2}\bigl(\ell^{\infty}/c_{00}\bigl)=I_{2}\bigl(c_{0}/c_{00}\bigl)=\{0\}$ and $\lambda\bigl(\ell^{\infty}/c_{00}\bigl)=\lambda\bigl(c_{0}/c_{00}\bigl)=2$.
  3. -
    $0\leqslant\lambda(X)<\omega_{1}$ for every POVS $X$.
  4. -
    $\lambda(X)=3$ for the vector lattice $X$ in Nakayama’s example.
We conclude with some (still open) questions related to Definitions 3.3 and 3.4.
$i)$ Is $\lambda(X)=\alpha(X)+1$ for every POVS $X$?
ii)ii) Is there a POVS $X$ with $\alpha(X)=\alpha$ (with $\lambda(X)=\alpha$) for each $\alpha\in\omega_{1}$?
$iii)$ Is there a vector lattice $X$ with $\alpha(X)=\alpha$ (with $\lambda(X)=\alpha$) for each $\alpha\in\omega_{1}$?

### v3

@@
 Abstract
-It is proved that, for a pre-ordered vector space $X$, the quotient space $(X/A,[W])$ is the Archimedeanization of $X$, where $W$ is the closure of the positive wedge $X_{+}$ in the ru-topology, $A=W\cap(-W)$, and $[W]$ is the quotient set of $W$ in $X/A$.
+It is proved that, for a pre-ordered vector space $X$, the quotient space $(X/A,[W])$ is the Archimedeanization of $X$, where $W$ is the closure of the positive wedge $X_{+}$ in the ru-topology, $A=W\cap-W$, and $[W]$ is the quotient set of $W$ in $X/A$.
 Eduard Emelyanov1
@@
 The present paper is devoted to an interplay between the Archimedean property and relatively uniform convergence in (pre-)ordered vector spaces.
-In Theorem 3.3, we prove that the quotient set $[W]$ of the $\tau_{ru}$-closure of the positive wedge $X_{+}$ of a pre-ordered vector space $X$ is an Archimedean cone in quotient space $X/(W\cap(-W))$.
-The main result of our paper is Theorem 3.5 which establishes that $(X/(W\cap(-W)),[W])$ is the Archimedeanization of $X$.
+In Theorem 3.3, we prove that the quotient set $[W]$ of the $\tau_{ru}$-closure of the positive wedge $X_{+}$ of a pre-ordered vector space $X$ is an Archimedean cone in quotient space $X/(W\cap-W)$.
+The main result of our paper is Theorem 3.4 which establishes that $(X/(W\cap-W),[W])$ is the Archimedeanization of $X$.
 2 Preliminaries
@@
 A subset $W$ of a vector space $X$ is a wedge if $W+W\subseteq W\ \ \text{\rm and}\ \ tW\subseteq W\ \ \text{\rm for}\ \ t\geqslant 0$.
-A wedge $C$ is a cone whenever $C\cap(-C)=\{0\}$.
+A wedge $C$ is a cone whenever $C\cap-C=\{0\}$.
 Under an ordered vector space (briefly, OVS) we understand a vector space $X$ together with a cone (denoted by $X_{+}$ and called a positive cone of $X$).
 Replacing a cone by a wedge, we also consider a pre-ordered vector space (briefly, POVS).
-A wedge $W$ in a vector space $X$ is majorizing (or generating) if
-[…]
-for every $x\in X$.
+A wedge $W$ in a vector space $X$ is generating if WW=XW-W=X.
 Every OVS (POVS) $X$ is equipped with a partial (pre-)order
 $x\leqslant y\Longleftrightarrow y-x\in X_{+}.$
-Every pair of vectors $\{a,b\}$ in an OVS (POVS) $X$ produces an order interval (possibly empty, if $a\not\leqslant b$):
+Each pair of vectors $\{a,b\}$ in an OVS (POVS) $X$ produces an order interval (possibly empty, if $a\not\leqslant b$):
 $a,b]:=\{x\in X:a\leqslant x\leqslant b\}=(a+X_{+})\cap(b-X_{+}).$
@@
     almost Archimedean if $\bigcap\limits_{n\in\mathbb{N}}\bigl[-\frac{1}{n}u,\frac{1}{n}u\bigl]=\{0\}$ for every $u\in X_{+}$.
-Clearly, each almost Archimedean POVS must be an OVS.
-Moreover, every Archimedean OVS is almost Archimedean.
-The converse is not true even for two-dimensional OVSs (see Example 2.5 in [3]).
-It is also straightforward to see: a POVS $X$ is almost Archimedean iff $X_{+}$ does not contain a straight line.
-A net $(x_{\alpha})$ in a POVS $X$ ru-converges to $x\in X$ (shortly, $x_{\alpha}\operatorname{\xrightarrow[]{{\it ru}}}x$) if, for some (a regulator of the convergence) $w\in X_{+}$, there exists a sequence $(\alpha_{n})$ of indices with $x_{\alpha}-x\leqslant\pm\frac{1}{n}w$ for $\alpha\geqslant\alpha_{n}$.
+Clearly, every almost Archimedean wedge is a cone and every Archimedean cone is almost Archimedean.
+On the other hand, even a two-dimensional almost Archimedean OVS need not to be Archimedean (see Example 2.5 in [3]).
+Furthermore, an OVS $X$ is almost Archimedean if and only if $X_{+}$ does not contain a straight line.
+A net $(x_{\alpha})$ in a POVS $X$ ru-converges to $x\in X$ (shortly, $x_{\alpha}\operatorname{\xrightarrow[]{{\it ru}}}x$) if, for some (regulator of the convergence) $w\in X_{+}$, there exists a sequence $(\alpha_{n})$ of indices with $\pm(x_{\alpha}-x)\leqslant\frac{1}{n}w$ for $\alpha\geqslant\alpha_{n}$.
 Whenever we need to specify a regulator $w$, we write $x_{\alpha}\operatorname{\xrightarrow[]{{\it ru}}}x(w)$.
@@
 The notion of ru-convergence was introduced by L.V.
-Kantorovich [9].
-3 Closedness of the positive wedge in ru-topology vs Archimedean property in pre-ordered vector spaces
-Let $X/A$ be a quotient space. A quotient set of $B\subseteq X$ is the set $B]=\{[x]\in X/A:x\in B\}$ of all equivalence classes formed from a set $B$ in the quotient space $X/A$.
+Kantorovich in [10].
+3 Closedness of a positive wedge in ru-topology vs Archimedean property
+Let $X/A$ be a quotient space.
+A quotient set of $B\subseteq X$ is the set $B]=\{[x]\in X/A:x\in B\}$ of all equivalence classes formed from a set $B$ in the quotient space $X/A$.
 The following lemma is a standard fact.
 Lemma 3.1.
-Let $W$ be a wedge in a vector space $X$. Then $[W]$ is a cone in $X/A$, where $A=W\cap(-W)$. Moreover, $[W]$ is majorizing in $X/A$ whenever $W$ is majorizing in $X$.
+Let $W$ be a wedge in a vector space $X$. Then $[W]$ is a cone in $X/A$, where $A=W\cap-W$. Moreover, $[W]$ is generating in $X/A$ if and only if $W$ is generating in $X$.
 Proof.
@@
 Then $x=x_{1}+A=-x_{2}+A$ for some $x_{1},x_{2}\in W$.
-It follows $x\in(W+W)\cap(-W-W)\subseteq W\cap(-W)=A$, and hence $[x]=0$.
-The rest is trivial.
+It follows $x\in(W+W)\cap(-W-W)\subseteq W\cap-W=A$, and hence $[x]=0$.
+Suppose WW=XW-W=X and let $x]\in X/A$. Then, $x=y_{1}-y_{2}$ for some $y_{1},y_{2}\in W$, and hence $x]=[y_{1}-y_{2}]=[y_{1}]-[y_{2}]\in[W]-[W]$. Thus, $[W]$ is generating in $X/A$.
+Suppose [W][W]=X/A[W]-[W]=X/A and let $x\in X$.
+Take $y_{1},y_{2}\in W$ with $x]=[y_{1}]-[y_{2}]=[y_{1}-y_{2}]$.
+So, for some $z\in A$ we have $x+z=y_{1}-y_{2}$, and hence $x=y_{1}-(y_{2}+z)\in W-W$.
+Therefore, $W$ is generating in $X$.
 ∎
-An OVS $X$ is a vector lattice if $x^{+}:=\inf\bigl\{X_{+}\cap(x+X_{+})\bigl\}$ exists for every $x\in X$.
+An OVS $X$ is a vector lattice if $x^{+}:=\sup\{x,0\}$ exists for every $x\in X$.
 It is easy to see that every almost Archimedean vector lattice is Archimedean.
@@
 Moore proved in [11] that the ru-closed subsets of a vector lattice $X$ are exactly the closed sets of the so-called ru-topology on $X$.
-This topology can be also introduced on an arbitrary POVS (cf., [8]).
+Such a topology can be also introduced on an arbitrary POVS [9].
 Definition 3.1.
@@
 It follows immediately from Definition 3.1 that the ru-topology on a POVS $X$ is the strongest topology $\tau$ on $X$ with the property $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$.
-The ru-topology is not linear in general.
-It is proved in [10, Theorem 4.2] that every lattice homomorphism between vector lattices is continuous in the ru-topology.
-Indeed, it is true for arbitrary positive operators between POVSs, and if additionally the positive wedge in the co-domain is majorizing, it is also true for every order bounded operator.
+Clearly, $\tau_{ru}$ is translation invariant.
+In general, the topology $\tau_{ru}$ is not linear.
+It is useful to describe constructively the $\tau_{ru}$-closure of a subset of a POVS.
+The idea is coming from [11], where the pseudo uniform closure of a subset of a vector lattice was studied.
+Adopting the idea to the POVS setting, for a subset $S$ of a POVS $X$ we denote $S^{(0)}_{ru}:=S$; $S^{(1)}_{ru}:=S^{\prime}_{ru}$ the set of all $x\in X$ such that $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ for some sequence $(x_{n})_{n}$ in $S$; $S^{(\alpha)}_{ru}:=\bigl(S^{(\alpha-1)}_{ru}\bigl)^{\prime}_{ru}$ when the ordinal $\alpha$ is not limit; and $S^{(\alpha)}_{ru}:=\bigcup\limits_{\beta\in\alpha}S^{(\beta)}_{ru}$ when $\alpha$ is limit.
+Lemma 3.2.
+Let $S$ be a subset of a POVS $X$. Then:
+  1. $i)$
+    $S^{(\alpha_{1})}_{ru}\subseteq S^{(\alpha_{2})}_{ru}$ whenever $\alpha_{1}\leqslant\alpha_{2}$.
+  2. ii)ii)
+    $\overline{S}_{ru}=\bigcup\limits_{\alpha\in\omega_{1}}S^{(\alpha)}_{ru}$, where $\omega_{1}$ is the first uncountable ordinal.
+  3. $iii)$
+    $S\subseteq X_{+}-X_{+}\Longleftrightarrow\overline{S}_{ru}\subseteq X_{+}-X_{+}$.
+Proof.
+$i)$ It is straightforward.
+ii)ii) It is similar to the proof of [11, Theorem 3.3], where the vector lattice case is considered.
+$iii)$ It suffices to prove $S\subseteq X_{+}-X_{+}\Longrightarrow S^{\prime}_{ru}\subseteq X_{+}-X_{+}$.
+So, let $S\subseteq X_{+}-X_{+}$ and $S\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$.
+By passing to a subsequence, if necessary, we may assume that $\pm(x_{n}-x)\leqslant\frac{1}{n}w$ for some $w\in X_{+}$ and all $n\in\mathbb{N}$.
+In particular, $x-x_{1}\leqslant w$, or $x\leqslant x_{1}+w$.
+Since $x_{1}=x_{1}^{1}-x_{1}^{2}$ for some $x_{1}^{1},x_{1}^{2}\in X_{+}$, we obtain $x\leqslant w+x_{1}^{1}$ and $x=(w+x_{1}^{1})-(w+x_{1}^{1}-x)\in X_{+}-X_{+}$.
+∎
+Lemma 3.3.
+Let $D$ be a wedge in a POVS $X$. Then:
+  1. $i)$
+    $D^{(\alpha)}_{ru}$ is a wedge in $X$ for every ordinal $\alpha$.
+  2. ii)ii)
+    $\overline{D}_{ru}$ is a wedge in $X$.
+Proof.
+$i)$ Let $\alpha$ be an ordinal.
+We have to prove that $t\cdot D^{(\alpha)}_{ru}\subseteq D^{(\alpha)}_{ru}$ for all $t\in\mathbb{R}_{+}$, and that $D^{(\alpha)}_{ru}+D^{(\alpha)}_{ru}\subseteq D^{(\alpha)}_{ru}$.
+By definition, $D^{(0)}_{ru}=D$, so $D^{(0)}_{ru}$ is a wedge.
+Suppose $D^{(\beta)}_{ru}$ is a wedge for every $\beta<\alpha$.
+First, consider the case when $\alpha=\beta+1$ for some $\beta$.
+So, $D^{(\beta)}_{ru}$ is a wedge and $D^{(\alpha)}_{ru}=\bigl(D^{(\beta)}_{ru}\bigl)^{\prime}_{ru}$.
+Let $t\geqslant 0$ and $x\in D^{(\alpha)}_{ru}$.
+Find a sequence $(x_{n})$ in $D^{(\beta)}_{ru}$ with $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$.
+By passing to a subsequence, we may assume that $\pm(x_{n}-x)\leqslant\frac{1}{n}w$ for some $w\in X_{+}$ and all $n\in\mathbb{N}$.
+Then, $\pm(tx_{n}-tx)\leqslant\frac{1}{n}tw$ for all $n\in\mathbb{N}$.
+We conclude $D^{(\beta)}_{ru}\ni tx_{n}\operatorname{\xrightarrow[]{{\it ru}}}tx$, and hence $tx\in\bigl(D^{(\beta)}_{ru}\bigl)^{\prime}_{ru}=D^{(\alpha)}_{ru}$.
+Now, let $\alpha$ be a limit ordinal, $t\geqslant 0$, and $x\in D^{(\alpha)}_{ru}=\bigcup\limits_{\beta\in\alpha}D^{(\beta)}_{ru}$.
+Then, $x\in D^{(\beta)}_{ru}$ for some $\beta<\alpha$.
+By the assumption, $tx\in tD^{(\beta)}_{ru}\subseteq D^{(\beta)}_{ru}\subseteq D^{(\alpha)}_{ru}$.
+We conclude again $tx\in D^{(\alpha)}_{ru}$.
+The proof of $D^{(\alpha)}_{ru}+D^{(\alpha)}_{ru}\subseteq D^{(\alpha)}_{ru}$ is similar, we point out only the case of $\alpha=\beta+1$.
+So, let $x,y\in D^{(\beta+1)}_{ru}$, say $D^{(\beta)}_{ru}\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ and $D^{(\beta)}_{ru}\ni y_{n}\operatorname{\xrightarrow[]{{\it ru}}}y$.
+By passing to subsequences, we may assume $\pm(x_{n}-x)\leqslant\frac{1}{n}w$ and $\pm(y_{n}-y)\leqslant\frac{1}{n}v$ for some $w,v\in X_{+}$ and all $n\in\mathbb{N}$.
+Then, $\pm((x_{n}+y_{n})-(x+y))\leqslant\frac{1}{n}(w+v)$ for all $n\in\mathbb{N}$.
+As $D^{(\beta)}_{ru}$ is a wedge, we conclude $D^{(\beta)}_{ru}\ni(x_{n}+y_{n})\operatorname{\xrightarrow[]{{\it ru}}}(x+y)$, and hence $x+y\in\bigl(D^{(\beta)}_{ru}\bigl)^{\prime}_{ru}=D^{(\beta+1)}_{ru}$.
+ii)ii) It follows from $i)$ due to Lemma 3.2. ∎
+It is proved in [11, Theorem 4.2] that every lattice homomorphism between vector lattices is continuous in the ru-topology.
+Indeed, it is true for arbitrary positive operators between POVSs, and if additionally the positive wedge in the co-domain is generating, it is also true for every order bounded operator.
 Lemma 3.4.
@@
   2. ii)ii)
-    If $Y_{+}$ is majorizing and $T$ is order bounded then $T$ is $\tau_{ru}$-continuous.
+    If $Y_{+}$ is generating and $T$ is order bounded then $T$ is $\tau_{ru}$-continuous.
 Proof.
@@
 To this end, let $T^{-1}(S)\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$.
-By passing to a subsequence, we may suppose $x_{n}-x\leqslant\pm\frac{1}{n}u$ for some $u\in X_{+}$ and all $n\in\mathbb{N}$.
+By passing to a subsequence, we may suppose $\pm(x_{n}-x)\leqslant\frac{1}{n}u$ for some $u\in X_{+}$ and all $n\in\mathbb{N}$.
 By the definition of $\tau_{ru}$-topology, it remains to show $x\in T^{-1}(S)$.
-If $T\geqslant 0$ then $Tx_{n}-Tx\leqslant\pm\frac{1}{n}Tu$, or $Tx_{n}\operatorname{\xrightarrow[]{{\it ru}}}Tx$.
+If $T\geqslant 0$ then $\pm(Tx_{n}-Tx)\leqslant\frac{1}{n}Tu$, or $Tx_{n}\operatorname{\xrightarrow[]{{\it ru}}}Tx$.
 Therefore, $S\ni Tx_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}Tx$ and, since $S$ is $\tau_{ru}$-closed, we infer $Tx\in S$, and hence $x\in T^{-1}(S)$.
-If $Y_{+}$ is majorizing and $T$ is order bounded, then $T$ is ru-continuous by [5, Theorem 2.4 ii)] As $T^{-1}(S)\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ then $S\ni Tx_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}Tx$.
-Thus, $Tx\in S$, and again $x\in T^{-1}(S)$.
+If $Y_{+}$ is generating and $T$ is order bounded, then $T$ is ru-continuous by [5, Theorem 2.4 ii)].
+Therefore, $T^{-1}(S)\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ implies $S\ni Tx_{n}\operatorname{\xrightarrow[]{{\it ru}}}Tx$.
+Thus, $S\ni Tx_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}Tx$.
+So, $Tx\in S$, and again $x\in T^{-1}(S)$.
 ∎
-It was observed by T. Ito [6] that a vector lattice $X$ is Archimedean iff $X_{+}$ is ru-closed. We show that it is happened also with every POVS.
-Lemma 3.3.
-A POVS $X$ with a positive wedge $X_{+}$ is Archimedean iff $X_{+}$ is ru-closed.
+It was observed by T. Ito [7] that a vector lattice $X$ is Archimedean iff $X_{+}$ is ru-closed. We show that it is happened also in every POVS.
+Proposition 3.1.
+A POVS $X$ is Archimedean if and only if its positive wedge $X_{+}$ is ru-closed.
 Proof.
@@
 […]
-Lemma 3.3 should be compared with the following classical fact (cf., [1, Lemma 2.3]).
+Proposition 3.1 should be compared with the following classical fact (cf., [1, Lemma 2.3]).
+Assertion 3.1.
+Every POVS $X$ that admits a linear topology $\tau$ for which $X_{+}$ is $\tau$-closed is Archimedean.
+It is worth mentioning a recent paper [6] by A.E.
+Gutman and I.A.
+Emelianenkov, where an exhaustive description of the class of locally convex spaces in which all Archimedean cones are topologically closed is obtained.
+However, the $\tau_{ru}$-topology is not necessarily linear.
+But, as it was mentioned above, $\tau_{ru}$-topology on a POVS $X$ is the strongest topology $\tau$ on $X$ such that $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$.
+This leads to the following consequence of Proposition 3.1.
 Proposition 3.2.
-If a POVS $X$ admits a linear topology $\tau$ for which its wedge $X_{+}$ is $\tau$-closed, then $X$ is Archimedean.
-The $\tau_{ru}$-topology is not necessarily linear.
-However, as it was mentioned above, the $\tau_{ru}$-topology on a POVS $X$ is the strongest topology $\tau$ on $X$ such that $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$.
-Therefore, we obtain the following corollary of Lemma 3.3.
-Corollary 3.1.
-A POVS $X$ is Archimedean iff $X_{+}$ is $\tau$-closed in some topology $\tau$ on $X$ possessing the property: $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$.
+Let $X$ be a POVS. The following assertions are equivalent.
+  1. $i)$
+    $X$ is Archimedean.
+  2. ii)ii)
+    $X_{+}$ is ru-closed.
+  3. $iii)$
+    $X_{+}$ is closed in the topology $\tau_{ru}$.
+  4. $iv)$
+    $X_{+}$ is closed in some topology $\tau$ on $X$ possessing the property: $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$.
+Proof.
+Equivalence $i)\Longleftrightarrow ii)$ is the statement of Proposition 3.1, $ii)\Longrightarrow iii)$ follows from Definition 3.1, and $iii)\Longrightarrow iv)$ is trivial.
+$iv)\Longrightarrow ii)$ Let $X_{+}$ be closed in a topology $\tau$ on $X$ satisfying $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$.
+If $X_{+}$ is not ru-closed, for some sequence $(x_{n})$ we have $X_{+}\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\notin X_{+}$, and hence $X_{+}\ni x_{n}\operatorname{\xrightarrow[]{\tau}}x$.
+Since $X_{+}$ is $\tau$-closed, we obtain $x\in X_{+}$, a contradiction.
+∎
 One of the most exciting results concerning interrelations between Archimedean cones and ru-convergence is the Luxemburg-Moore-Veksler theorem:
-Theorem 3.2.
+Assertion 3.2.
 ([11, 14]) Let $A$ be a lattice ideal in a vector lattice $X$. Then
 […]
-The forward implication of Theorem 3.2 is almost trivial in the OVS case [3, Proposition 3.2(a)].
-We do not have yet an extension of Theorem 3.2 to any reasonable class of POVS beside vector lattices.
+The forward implication of Assertion 3.2 is almost trivial in the OVS case [3, Proposition 3.2(a)].
+On the other hand, for an OVS $V$ from [3, Example 2.5], $V=V/\{0\}$ is not Archimedean although the ideal $\{0\}$ is ru-closed.
+We do not have yet an extension of Assertion 3.2 to reasonable classes of POVSs beside vector lattices.
 However, for a very natural ru-closed order ideal $A$ in an arbitrary POVS $X$ the quotient $X/A$ turns to be Archimedean.
@@
 Let $X$ be a POVS with a positive wedge $X_{+}$.
-Then $[W]$ is an Archimedean cone in $X/A$, where $W$ is the $\tau_{ru}$-closure of $X_{+}$ and $A=W\cap(-W)$.
-Moreover, the cone $[W]$ is majorizing whenever $X_{+}$ is majorizing.
+Then $[W]$ is an Archimedean cone in $X/A$, where $W$ is the $\tau_{ru}$-closure of $X_{+}$ and $A=W\cap-W$.
+Moreover, the cone $[W]$ is generating whenever $X_{+}$ is generating.
 Proof.
-Clearly, $A$ is an order ideal in $X$.
-By Lemma 3.1, $[W]$ is a cone in $X/A$.
+By Lemma 3.3, $W$ is a wedge.
+Lemma 3.1 implies that $[W]$ is a cone in $X/A$.
 The set $W$ (and, also W-W) is $\tau_{ru}$-closed.
@@
 So, there exists $(a_{n})_{n}$ in $A$ such that $x-\frac{1}{n}y+a_{n}\in-W$ for all $n\in\mathbb{N}$.
-Note that $a_{n}\operatorname{\xrightarrow[]{{\it ru}}}0$ (with a regulator 0).
-As W-W is $\tau_{ru}$-closed, we infer from $\frac{1}{n}y\operatorname{\xrightarrow[]{{\it ru}}}0$ and $a_{n}\operatorname{\xrightarrow[]{{\it ru}}}0$ that $x-\frac{1}{n}y+a_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}x$, and hence $x\in-W$.
+Clearly, $a_{n}\operatorname{\xrightarrow[]{{\it ru}}}0$ (with regulator 0).
+By Lemma 3.2, $y\in X_{+}-X_{+}$.
+Then, $\pm y\leqslant u$ for some $u\in X_{+}$, and hence $\frac{1}{n}y\operatorname{\xrightarrow[]{{\it ru}}}0$ (with regulator $u$).
+So, $x-\frac{1}{n}y\operatorname{\xrightarrow[]{{\it ru}}}x$.
+As W-W is $\tau_{ru}$-closed, we infer from $x-\frac{1}{n}y\operatorname{\xrightarrow[]{{\it ru}}}x$ and $a_{n}\operatorname{\xrightarrow[]{{\it ru}}}0$ that $x-\frac{1}{n}y+a_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}x$, and hence $x\in-W$.
 It follows $x]\in-[W]$, or $x]\leqslant 0$.
-So, $(X/A,[W])$ is Archimedean.
-Due to Lemma 3.1, $[W]$ is majorizing whenever $X_{+}\subseteq W$ is majorizing.
-∎
+Therefore, $(X/A,[W])$ is Archimedean.
+Suppose $X_{+}$ is generating. Since $X_{+}\subseteq W$ then $W$ is generating too, and hence $[W]$ is generating by Lemma 3.1. ∎
 It was proved in [13] that every OVS with an order unit has a unique (up to an order isomorphism) Archimedization.
-This result was extend further to arbitrary OVS in [3].
-The following definition was introduced in [3, Definition 5].
+This result has being extend to arbitrary OVS in [3].
+The next notion was introduced in [3, Definition 5].
 Definition 3.2.
@@
 If $\left\langle Y_{0},\phi_{0}\right\rangle$ is an initial object of $\mathcal{C}_{Arch}(X)$, then the OVS $Y_{0}$ is called the Archimedeanization of $X$.
-By [3, Corollary 3.5], any almost Archimedean cone can be embedded into an Archimedean cone.
-In particular, the Archimedeanization of any almost Archimedean OVS space $X$ can be obtained via replacing $X_{+}$ by the intersection of all Archimedean cones containing $X_{+}$.
-The following result is proved in [3, Theorem 1].
-Proposition 3.4.
-Every OVS has the Archimedeanization.
-We present a construction of the Archimedeanization (based on Theorem 3.3), which seems to be more elementary than one given in [3].
+The following result is proved in [3].
+Assertion 3.3.
+([3, Theorem 1]) Every OVS has the Archimedeanization.
+We present a construction of the Archimedeanization (based on Theorem 3.3), which looks more elementary than the one given in [3].
 Theorem 3.4.
-Let $X$ be a POVS with a positive wedge $X_{+}$. Let $W$ is the $\tau_{ru}$-closure of $X_{+}$ and $A=W\cap(-W)$. Then $(X/A,[W])$ is the Archimedeanization of $X$.
+Let $X$ be a POVS with a positive wedge $X_{+}$, $W$ the $\tau_{ru}$-closure of $X_{+}$, and $A=W\cap-W$. Then $(X/A,[W])$ is the Archimedeanization of $X$.
 Proof.
@@
 It remains to find a positive operator $T:(X/A,[W])\to Y$ such that $T\circ\phi_{0}=\phi$.
-Since $Y$ is Archimedean, it follows from Lemma 3.3 that $Y_{+}$ is ru-closed, and hence is $\tau_{ru}$-closed.
+Since $Y$ is Archimedean, it follows from Proposition 3.1 that $Y_{+}$ is ru-closed, and hence is $\tau_{ru}$-closed.
 First, we show $\phi(W)\subseteq Y_{+}$.
@@
 As $Y_{+}$ is $\tau_{ru}$-closed, $\phi(x)\in Y_{+}$.
-Thus, $\phi(W)\subseteq Y_{+}$ and $\phi(-W)\subseteq-Y_{+}$, and hence $\phi(A)\subseteq Y_{+}\bigcap(-Y_{+})=\{0\}$.
+Thus, $\phi(W)\subseteq Y_{+}$ and $\phi(-W)\subseteq-Y_{+}$, and hence $\phi(A)\subseteq Y_{+}\bigcap-Y_{+}=\{0\}$.
 So, $T:(X/A,[W])\to Y$ such that $T[x]:=\phi(x)$ is well defined and positive.
@@
 ∎
-For applications of Theorem 3.5, it is useful to describe constructively the $\tau_{ru}$-closure of the positive wedge $X_{+}$ of a POVS $X$.
-The idea comes from [10], where the pseudo uniform closure of a subset of a vector lattice was studied.
-Adopting the idea to the POVS setting, for a subset $S$ of a POVS $X$ we denote by $S^{\prime}_{ru}$ the set of all $x\in X$ such that $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ for some sequence $(x_{n})_{n}$ in $S$.
-Denote by $S^{(\alpha)}_{ru}$ the set $S^{(\alpha-1)}_{ru}$ when the ordinal $\alpha$ is not limit, otherwise $S^{(\alpha)}_{ru}:=\bigcup\limits_{\beta\in\alpha}S^{(\beta)}_{ru}$.
-It is straightforward to see that $S^{(\alpha_{1})}_{ru}\subseteq S^{(\alpha_{2})}_{ru}$ whenever $\alpha_{1}\leqslant\alpha_{2}$, and $\overline{S}_{ru}=\bigcup\limits_{\alpha\in\omega_{1}}S^{(\alpha)}_{ru}$, where $\omega_{1}$ is the first uncountable ordinal (cf., [10, Theorem 3.3] in the vector lattice setting).
-In particular, we have $\overline{S}_{ru}=S^{\prime}_{ru}\ \Longleftrightarrow\ S^{\prime}_{ru}=(S^{\prime}_{ru})^{\prime}_{ru}$.
-So, in the notations of Theorem 3.5, $W=\overline{(X_{+})}_{ru}=\bigcup\limits_{\alpha\in\omega_{1}}(X_{+})^{(\alpha)}_{ru}$.
-Note that each $(X_{+})^{(\alpha)}_{ru}$ is a wedge in $X$.
-The statement of Theorem 3.5 motivates the following notion.
+Lemma 3.3 implies (in the notations of Theorem 3.4) that $W=\overline{(X_{+})}_{ru}=(X_{+})^{(\omega_{1})}_{ru}$. Theorem 3.4 motivates the following.
 Definition 3.3.
 A POVS $X$ has the Archimedeanization type $\alpha(X)$, whenever $\alpha(X)$ is the first ordinal $\alpha$ such that $\overline{(X_{+})}_{ru}=(X_{+})^{(\alpha)}_{ru}$.
-It can be seen easily:
+Let $X$ be a POVS.
+Following [3], denote $D_{X}:=\{x\in X|(\exists\xi\in X_{+})(\forall n\in\mathbb{N})\ x+n^{-1}\xi\in X_{+}\}$, and $N_{X}:=\bigl\{x\in X:(\exists y\in X)(\forall n\in\mathbb{N})\pm x\leqslant\frac{1}{n}y\bigl\}$ the set of all infinitesimals of $X$.
+For every $x,y\in X$ denote $\Omega^{x}_{y}:=\{\alpha\in\mathbb{R}:x+\alpha y\in X_{+}\}$.
+Lemma 3.5.
+Let $X$ be a POVS and $y\in X$. The following assertions are equivalent:
+  1. $i)$
+    $\Omega^{x}_{y}\neq\{0\}$ for some $x\in X_{+}$.
+  2. ii)ii)
+    $y\in X_{+}-X_{+}$.
+Proof.
+$i)\Longrightarrow ii)$ Let $x\in X_{+}$ satisfy $\Omega^{x}_{y}\neq\{0\}$.
+Then $x+\alpha y\in X_{+}$ for some $\alpha\neq 0$.
+Thus, $\alpha y=(x+\alpha y)-x\in X_{+}-X_{+}$, and hence $y=\alpha^{-1}(x+\alpha y)-\alpha^{-1}x\in X_{+}-X_{+}$.
+$ii)\Longrightarrow i)$ Let $y\in X_{+}-X_{+}$.
+Then $y=x^{\prime}-x^{\prime\prime}$ for some $x^{\prime},x^{\prime\prime}\in X_{+}$.
+It follows $y\geqslant-x^{\prime\prime}$, and $x^{\prime\prime}+\alpha y\geqslant x^{\prime\prime}-\alpha x^{\prime\prime}\geqslant 0$ for all $\alpha\in[0,1]$, and hence $0,1]\subseteq\Omega^{x^{\prime\prime}}_{y}$.
+∎
+Obviously, $X_{+}\subseteq D_{X}=\{x\in X|(\exists\xi\in X_{+})(\forall r>0)\ x+r\xi\in X_{+}\}$.
+In particular, for every $x\in D_{X}$ there exists $\xi\in X_{+}$ such that $\mathbb{R}_{+}\subseteq\Omega^{x}_{\xi}$.
+It is clear that $D_{X}$ is a wedge.
+Proposition 3.5.
+Let $X$ be a POVS. Then the following holds:
+  1. $i)$
+    $D_{X}=(X_{+})^{(1)}_{ru}$
+  2. ii)ii)
+    $D_{X}\cap-D_{X}=N_{X}$.
+  3. $iii)$
+    If $X$ is almost Archimedean then $D_{X}$ is a cone.
+Proof.
+$i)$ If $x\in D_{X}$ then, for some $\xi\in X_{+}$ we have $x+n^{-1}\xi\in X_{+}$ for all $n\in\mathbb{N}$.
+Since $x+n^{-1}\xi\operatorname{\xrightarrow[]{{\it ru}}}x$, then $x\in(X_{+})^{(1)}_{ru}$.
+Thus, $D_{X}\subseteq(X_{+})^{(1)}_{ru}$.
+If $x\in(X_{+})^{(1)}_{ru}$, there exist $\xi\in X_{+}$ and a sequence $(x_{n})$ in $X_{+}$ satisfying $\pm(x_{n}-x)\leqslant n^{-1}\xi$ for all $n\in\mathbb{N}$.
+Then $0\leqslant x_{n}\leqslant x+n^{-1}\xi$, and hence $x+n^{-1}\xi\in X_{+}$ for all $n\in\mathbb{N}$.
+So, $x\in D_{X}$.
+Thus, $(X_{+})^{(1)}_{ru}\subseteq D_{X}$.
+ii)ii) Clearly, $N_{X}$ is an order ideal in $X$ and $N_{X}\subseteq(X_{+})^{(1)}_{ru}$. Thus, $i)$ implies $N_{X}\subseteq D_{X}\cap-D_{X}$.
+Suppose $x\in D_{X}\cap-D_{X}$. Then, for some $\xi_{1},\xi_{2}\in X_{+}$ we have
+$x+r\xi_{1}\subseteq X_{+}\ \ \text{\rm and}\ \ -x+r\xi_{2}\subseteq X_{+}\ \ \ \ (\forall r>0).$
+It follows $r(\xi_{1}+\xi_{2})\leqslant x\leqslant r(\xi_{1}+\xi_{2})$ for all $r>0$, and hence $x\in N_{X}$. So, $D_{X}\cap-D_{X}\subseteq N_{X}$.
+$iii)$ Since $X$ is almost Archimedean then $N_{X}=\{0\}$, and hence $D_{X}\cap-D_{X}=\{0\}$ by ii)ii). ∎
+In general, $D_{X}$ is not necessarily a cone even if $\text{\rm dim}(X)=2$ (take, for example, $\mathbb{R}^{2}_{lex}$).
+It can be seen:
   1. -
@@
     $0\leqslant\alpha(X)<\omega_{1}$ for every POVS $X$, and $\alpha(X)=2$ for the vector lattice $X$ in Nakayama’s example (cf., Example 6.4 in [11]).
-Let us consider the following order ideals in a POVS $X$: the set $I_{0}(X):=\{0\}$ of all infinitesimals of $X$ of order 0; the set:
-$I_{1}(X)=I(X):=\bigl\{x\in X:(\exists y\in X)(\forall n\in\mathbb{N})\pm x\leqslant\frac{1}{n}y\bigl\}$
-of all infinitesimals of order 1 (or, just infinitesimals of $X$); and the set
+Let us consider the following order ideals in a POVS $X$: the set $I_{0}(X):=\{0\}$ of all infinitesimals of $X$ of order 0; the set: $I_{1}(X):=N_{X}$ of all infinitesimals of $X$); and the set
 $I_{\alpha}(X):=\bigl\{x\in X:[x]\in I\bigl(X/_{\bigcup_{\gamma\in\alpha}I_{\gamma}(X)}\bigl)\bigl\}.$
@@
 It can be seen that $I_{\alpha_{1}}(X)\subseteq I_{\alpha_{2}}(X)\subseteq(X_{+})^{(\alpha_{2})}_{ru}$ for $\alpha_{1}\leqslant\alpha_{2}$.
-Clearly, a POVS $X$ is almost Archimedean iff $I(X)=\{0\}$.
-If a POVS $X$ has a nonzero order unit $u\in X$ (i.e., $u$ satisfies $X=\bigcup\limits_{n\in\mathbb{N}}[-nu,nu]$) then $u\notin I(X)$.
+If a POVS $X$ has a nonzero order unit $u\in X$ (i.e., $u$ satisfies $X=\bigcup\limits_{n\in\mathbb{N}}[-nu,nu]$) then $u\notin N_{X}$.
 Definition 3.4.